A man waters, with a water hose, with an area of 2.26'. If the volume of the water coming out of the hose fills a 10 quart pail in 2 minutes, what is the rainfall depth (in inches ) equivalent if he runs the garden hose for 12 minutes? 1gallon = 8pounds 1 cubic foot = 62.4 pounds
1
Expert's answer
2016-02-10T00:01:18-0500
Solution:
The volume of water flowing out of the hose for 12минут equal
V = 10*12/2 = 60 quarts = 15 gallons
Translate the volume of water in cubic feet
V = 15*8/62.4 = 1.92 cubic foot
If the area, pour water equal to 2,26 square feet, the thickness of water equal to 1.92*12/2,26 = 10,21 inch.

Answer: Тhe rainfall depth equal 10,21 inch